OpenStudy (anonymous):

(logx)^3=log(x^3) please explain!! thank youuu!!

OpenStudy (anonymous):

i made logx=a sooooooo a^3-3a=0? a(a^2-3)=0 a=0 sooo x=1 a=±sqrt 3? is it right that i do ±? soooo it's 10^±sqrt 3 also? is there a way to simplify this further?

OpenStudy (anonymous):

@hartnn ? :D anyone?? :D

hartnn (hartnn):

10^(-sqrt3) = 1/ (10^(sqrt 3)) that the only thing i can think of.....

hartnn (hartnn):

and yes, x=1,10^±sqrt 3 is correct

hartnn (hartnn):

and yes it is right to do +/-

OpenStudy (anonymous):

yay okay thank youuu!!!:)

hartnn (hartnn):

welcome ^_^